Evaluate the given expression, leave the answer in simple radical form. sec60 degrees * tan135 degrees - tan60degrees * cot60
1
Expert's answer
2014-06-04T02:34:40-0400
Using the table meanings pf trigonometric functions we get
sec60 * tan135 - tan 60 * cot60=1/cos 60 * (-1) - \sqrt (3)*1/\sqrt (3)=-2-1=-3